4) Suppose that demand is given by Q = 1500 - 10P and supply by Q = 5P.

A) Find the equilibrium price and quantity. (6 pts.)

B) At the equilibrium price and quantity, find the price elasticity of demand and the price elasticity of supply. (6 pts.)
Answer #1

Answer

A)

equilibrium is at Qd=Qs

equating both equations

1500-10P=5P

15P=1500

P=100

Q=1500-10*100=500

the equilibrium price is $100 and the quantity is 500 units.

==========

B)

Р до Price elasticity of demand = *** до д(1500 — 10P) = –10 ӘР ӘР Р= 100 and Q = 500 100 E = -10 + son = -2

The elasticity of demand is -2.

=======

Price elasticity of supply = әр *о до Р де д (5Р) - 5 Р= 100 and Q = 500 E=5100 * 500 = 1

Price elasticity of supply is 1.